OpenStudy (anonymous):

Someone please! Help Me ! Find the area of a regular hexagon given the side length is 4 in. Leave your answer in exact form (with square root) and no decimals. (Use sq rt as the square root symbol.

OpenStudy (anonymous):

A regular hexagon is made up of 6 equilateral triangles, so 100 = 6(1/4 sqrt(3))* s^2 solve for s: 100 = 3/2 sqrt(3) s^2 s^2 = 100/(3/2 sqrt(3)) s^2 = 200/(3 sqrt(3)) s = sqrt(200/(3sqrt(3)) ~ = 6.2 cm

OpenStudy (ranga):



OpenStudy (ranga):

AB = 4 AC = AB/2 = 2 tan(60) = OC / AC sqrt(3) = OC / 2 OC = 2 * sqrt(3) Area of triangle = 1/2 * base * height = 1/2 * AB * OC = 1/2 * 4 * 2 * sqrt(3) = 4 * sqrt(3) Area of hexagon = 6 * area of triangle OAB = 6 * 4 * sqrt(3) = 24sqrt(3)

OpenStudy (anonymous):

thank you so much :) !!!!

OpenStudy (ranga):

you are welcome.